# Surface to Seafloor: A Generative AI Framework for Decoding the Ocean Interior State

### v2

@@ 1 Introduction
Several physically motivated methods have been developed to infer subsurface ocean states from surface data, each with distinct advantages and limitations. The Surface Quasi-Geostrophic (SQG) framework has been widely used for subsurface reconstruction. Still, it is constrained to scales above 150 km (Held et al., 1995; Lapeyre and Klein, 2006; Wang et al., 2013; Liu et al., 2019) due to the inherent assumptions in QG theory and the lack of detailed information about the vertical stratification. Recent satellite missions such as the Surface Water and Ocean Topography (SWOT) Mission (Fu et al., 2024) offer much higher resolution (down to 1–10 km), limiting the applicability of traditional large-scale geostrophic approximations. Theoretical extensions, such as the SQG-mEOF-R method, improve upon standard SQG by accurately reconstructing both surface- and subsurface-intensified eddies, particularly in regions with complex vertical stratification, making it more effective across oceanic regimes such as the Kuroshio Extension and the Peru-Chile upwelling system, (Yan et al., 2020), but still build on assumptions applicable only to larger scale ocean structures.
@@
While these studies have demonstrated that satellite-derived surface information can infer the ocean interior, most have relied on supervised learning, where a model is trained to minimize residuals against a single “true” state. However, a given surface ocean state does not deterministically correspond to a unique deep ocean state. A more physically sound approach requires probabilistic modeling, which allows for generating an ensemble of plausible subsurface states constrained by surface observations, while preserving the uncertainties and variability inherent in ocean dynamics.
Importantly, probabilistic modeling allows for uncertainty quantification (UQ) (Forget et al., 2015; Souza et al., 2020): not only can samples from the inferred distribution be explored, but these samples also provide estimates of plausible subsurface fields when surface information is insufficient to constrain the output fully. Generative modeling offers a flexible approach to sampling from such distributions, with score-based diffusion models emerging as a particularly effective tool for high-dimensional statistical inference. Generative-adversarial models (GANs) have also been applied to predict ocean temperature profiles from sea surface temperature (Meng et al., 2021; Zhang et al., 2023), but these works did not leverage the potential for uncertainty quantification. More recently, Martin et al. (2025) have leveraged score-based generative methods in the context of ocean surface data assimilation but again, they did not focus on uncertainty quantification. Beyond conditional sampling, score-based generative models have demonstrated utility in diverse applications, including calculating response functions (Giorgini et al., 2024, 2025b), estimating the dimensionality of data manifolds (Stanczuk et al., 2023), stabilizing the dynamical evolution neural network models (Pedersen et al., 2025), or devising stochastic analogs to chaotic equations (Giorgini et al., 2025a).
@@ 2.1 Ocean Simulation
As with any idealized simulation, we focus on particularly salient features that generalize to more complex scenarios. Thus, the role of the simulation is not to capture realistic features of the ocean but rather to serve as a testing ground for whether or not statistical inference of the free surface to the ocean interior can be performed. With this in mind, we choose a baroclinic double gyre, similar to (Cox and Bryan, 1984), as a representative ocean simulation for the methodology. The simulation generates circulation patterns consistent with those observed in the subtropical and subpolar gyres. We will not run the ocean model to statistical equilibrium so that we can study both statistically stationary and non-stationary features, (Hasselmann, 1976).
Using Oceananigans.jl (Ramadhan et al., 2020; Silvestri et al., 2024a; Wagner et al., 2025a; Silvestri et al., 2024b; Wagner et al., 2025b), we solve the Boussinesq equations under the hydrostatic approximation. The dynamical core uses a nonlinear free surface and $z^{\star}$ coordinates. The prognostic variables are the horizontal velocities, $U$ and $V$, the SSH $\eta$, and the temperature $T$. A linear equation of state relates the buoyancy of seawater $B$ to the temperature, (Roquet et al., 2015). The figures will refer to the Oceananigans simulation output as “OcS” for brevity.
The idealized ocean domain extends from 15 to 75 degrees north in latitude and from 0 to 60 degrees in longitude with walls on all sides. The domain is 1800 m deep with a flat bottom. A zonal wind stress is applied at the surface, independent of longitude, with a cosine profile peaking eastward (positive) in the middle of the domain at a value of 0.1 $0.1\text{ N}\text{m}^{-2}$ and at a minimum and westward (i.e. negative) at the northern and southern walls. The surface temperature is restored to a linear profile in latitude ranging from $30$ C on the southern boundary at $15^{\circ}$, to 00 C at the northern boundary at $75^{\circ}$. A quadratic bottom drag is imposed at the seafloor and convective adjustment is used for vertical mixing. The (dimensionless) quadratic drag coefficient is $10^{-3}$ and we use a horizontal diffusivity and viscosity of $2.5\times 10^{2}\text{ m}^{2}\text{s}^{-1}$, a background vertical diffusivity of $10^{-5}\text{ m}^{2}\text{s}^{-1}$, and background vertical viscosity of $10^{-2}\text{ m}^{2}\text{s}^{-1}$.
@@
Our simulation is not in statistical equilibrium, as can be observed from Figure 2, which shows the $L^{1}$ norm (the average absolute value) of all variables at a few selected depths, e.g.
@@
We use this measure as it is sensitive to localized fluctuations, such as the root-mean-squared value. Including metric terms does not change the overall trend and is unnecessary for the present case, but we will include them later when calculating physically meaningful quantities rather than statistical measures of non-stationarity. The initial spinup period (blue) is discarded. Thus, all the data in the work uses the time periods associated with the middle portion (red) and the end portion (orange) of the simulation. Closer to the surface (first column), it appears that a quasi-statistical equilibrium has been reached, whereas at depth (last column), statistical equilibrium has not been reached, especially for temperature. This quasi-equilibrium nature of the ocean is a feature we wished to capture in our dataset when investigating the limits of statistical inference predicated on past data. We comment that the sea surface height’s $L^{1}$ norm is also not in equilibrium. Stated differently, our “data distribution” (red) will be different than our “validation distribution” (orange) and will contribute as an extra source of error.
@@ 2.2 Score-Based Diffusion Model
In the present work, we draw samples from subsurface fields, $U,V,W,T$, given the sea surface height $\eta$, where the underlying probability distribution comes from the primitive equations. Mathematically, this is represented as a conditional probability distribution, $\mathcal{P}$,
[…]
The conditional probability distribution outputs four fields $U,V,W,T$ at every grid point in three-dimensional space, conditioned on the SSH $\eta$. Concretely, samples from this distribution produce $4\times 256\times 256\times 15$ numbers conditioned on $256\times 256$ numbers using the full resolution of the simulation. This high-dimensional conditional probability distribution is learned from the data set.
If the conditional information is perfectly informative and there is a functional relationship, $\mathcal{F}$, between SSH and predicted fields and predicted fields. In this case, the conditional distribution collapses to a delta function,
$\mathcal{P}(\text{subsurface fields}|\text{surface fields})=\delta(\text{subsurface fields}-\mathcal{F}[\text{subsurface fields}]).$ (3)
If the provided information is “not useful” or maximally uninformative, then the conditional probability distribution, $\mathcal{P}$, should reduce down to the marginal distribution of the data, e.g., a random sample with statistics similar those the generative model has been trained on independently of the surface information,
where $\mathcal{P}_{M}$ is the marginal distribution of the subsurface fields. Most inference tasks lie between these two extremes, Equations 3 and 4. We comment that we use the word “marginal” as is commonly used in the statistics literature with regards to distributions.
Traditional machine learning methods find a functional form $\mathcal{F}$ between input and outputs. This methodology works best when such a functional form exists. In the absence of an approximate functional relationship, traditional methods that minimize a root-mean square (RMS) error with the aim of finding a functional form can only infer the average value of the conditional probability in Equation 2. (This happens for the same reason that a “mean” statistic minimizes the RMS error of a random variable.) We surmise that this observation partially explains why fields drawn from a conditional distribution better represent the statistical properties of fluid flows compared with those learned via a deterministic functional relationship. Details will not be “washed away” from averaging.
@@ 3.1 Baseline Performance of Generative Model
We see an overall agreement between the ground truth (second column) and the expected value of the model predictions (third column). This result means that a substantial ageostrophic component is also captured. The ensemble mean of the generated samples looks similar to the ground truth and the representative generated samples since there is little diversity in the generated samples. The conditional distribution is nearly a delta function, suggesting that the sea surface height is very informative of the ocean state and that a functional relationship can be found between the sea surface height and sub-surface currents. In other words, the present inference task is leaning towards Equation 3.
The sixth column shows the absolute value of the difference between the OcS and the AI mean, and the seventh column shows the ensemble standard deviation of the AI method. We represent predictive uncertainty by using the generated fields’ standard deviation at each grid point. The largest standard deviation is associated with the northern convecting region, especially for vertical velocity. We surmise that the sea surface height alone is insufficient to constrain the pointwise vertical velocity magnitude in these regions. The second largest source of uncertainty is the eddying gyre region, which contributes to the uncertainty of the velocity components. The mean discrepancy column and the AI Std column have similar spatial structures, suggesting that the method can identify regions in which it is less capable of prediction.
@@ 3.2 Ocean State Predictions and Uncertainties with Degraded Surface Height Input
These observations suggest that an SSH coarse-grained beyond $\mathcal{O}(1,000)$ kilometers is no longer informative since the generated samples become so diverse as to reproduce statistics of the data distribution, a “worst case” scenario for the generative method. As we coarse-grain the free surface (lose information), the inference task moves from something informative (similar to Equation 3) towards the uninformative (similar to Equation 4) marginal distribution of the training data.
@@ 3.3 Heat Flux Predictions and Uncertainties with Degraded Surface Height Input
¯=$\overline{\cdot}\stackrel{{\scriptstyle\textrm{def}}}{{=}}\frac{1}{60^{\circ}}\int_{0}^{60^{\circ}}$ $\cdot d\lambda,$ (5)
@@
We also calculate the eddy heat fluxes as a function of SSH resolution at a $400$ meter depth in Figure 8. Similar to the previous plot, the top row is the vertical heat flux, the bottom row is the meridional heat flux, the “ground truth” results are the blue lines, and the generative AI results with 100 samples are shown as a ribbon with $0.6,0.7,0.8,0.9$ quantiles from darker to lighter shadings of red. The heat flux prediction grows in uncertainty as we coarse-grain the sea surface height. The loss of information here is understood upon examining the first column of Figure 4. As the sea surface height (SSH) field is increasingly coarse-grained, we lose precise information about the location of the jet separation point — the region where the western boundary current detaches from the wall and begins to meander eastward. This separation point is critical in organizing mesoscale eddy activity and lateral heat transport. When spatial detail in the SSH is removed, the model can no longer resolve where the jet begins to meander, leading to increased uncertainty in the predicted eddy heat fluxes, particularly in the downstream region where the current interacts with ambient stratification.

### v3

@@ 1 Introduction
Several physically motivated methods have been developed to infer subsurface ocean states from surface data, each with distinct advantages and limitations. The Surface Quasi-Geostrophic (SQG) framework has been used for subsurface reconstruction. The theory works well when the inherent assumptions in QG theory are satisfied and there exists information about the vertical stratification (Held et al., 1995; Lapeyre and Klein, 2006; Wang et al., 2013; Liu et al., 2019). Theoretical extensions, such as the SQG model with a multivariate Empirical Orthogonal Function have been shown to improve upon standard SQG by accurately reconstructing both surface- and subsurface-intensified eddies, particularly in regions with complex vertical stratification like the Kuroshio Extension and the Peru-Chile upwelling system (Yan et al., 2020). However, recent satellite missions such as the Surface Water and Ocean Topography (SWOT) Mission (Fu et al., 2024) achieve higher resolution than traditional altimeters (down to 10 km) and resolve scales where the the QG approximations are no longer valid, necessitating additional theoretical development or alternative approaches.
@@
While these studies have demonstrated that satellite-derived surface information can be used to infer the ocean interior, most have relied on supervised learning, where a model is trained to minimize residuals against a single “true” state. However, a given surface ocean state does not deterministically correspond to a unique deep ocean state. A more physically sound approach requires probabilistic modeling, which allows for generating an ensemble of plausible subsurface states constrained by surface observations, while preserving the uncertainties and variability inherent in ocean dynamics. Probabilistic methods can be considered an extension of supervised learning methods as they can often reproduce the inference from such methods, as a subset of their capabilities. For example, minimizing the root-mean-square error of a probabilistic output results in the mean of the underlying distribution which is the target of supervised learning methods (barring discrepancies in training and architecture choices).
Importantly, probabilistic modeling allows for uncertainty quantification (UQ) (Forget et al., 2015; Souza et al., 2020): not only can samples from the inferred distribution be explored, but these samples also provide estimates of plausible subsurface fields when surface information is insufficient to constrain the output fully. Generative modeling offers a flexible approach to sampling from such distributions, with score-based diffusion models emerging as a particularly effective tool for high-dimensional statistical inference. Generative-adversarial models (GANs) have also been applied to predict ocean temperature profiles from sea surface temperature (Meng et al., 2021; Zhang et al., 2023), but these works did not leverage the potential for uncertainty quantification. The study in Martin et al. (2025) leveraged score-based generative methods in the context of ocean surface data assimilation but again, they did not focus on uncertainty quantification; however, there have been recent efforts at examining uncertainty quantification in the presence of sparse ocean-like measurements in a quasi-geostrophic model Babu et al. (2025). Score-based generative methods in particular have also been explored when addressing extreme event sampling as well Stamatelopoulos and Sapsis (2025). Beyond conditional sampling, score-based generative models have demonstrated utility in diverse applications, including calculating response functions (Giorgini et al., 2024, 2025d), estimating the dimensionality of data manifolds (Stanczuk et al., 2023), stabilizing the dynamical evolution neural network models (Pedersen et al., 2025), parameter calibration (Giorgini et al., 2025c), data-assimilation (Manshausen et al., 2025), or devising stochastic analogs to chaotic equations (Giorgini et al., 2025a; Giorgini, 2025; Giorgini et al., 2025b).
@@ 2.1 Ocean Simulation
We illustrate the generative AI method with an idealized simulation rather than reanalysis data or a more realistic ocean setup for two primary reasons: 1) with an idealized study it is possible to generate (re-generate) as much training data as needed to test the methodology and 2) a more realistic ocean model introduces additional challenges like measurement uncertainty, data gaps, and nonstationary statistics. We believe that one must first demonstrate that the proposed methodology works in an idealized setting here before attempting it in a more complex scenario.
As with any idealized simulation, we focus on particularly salient features that generalize to more complex scenarios. Thus, the role of the simulation is not to capture realistic features of the ocean but rather to serve as a testing ground for whether or not statistical inference of the free surface to the ocean interior can be performed. With this in mind, we choose a baroclinic double gyre, similar to (Cox and Bryan, 1984), as a representative ocean simulation for the methodology. The simulation generates circulation patterns consistent with those observed in the subtropical and subpolar gyres. We will not run the ocean model to full statistical equilibrium so that we can study both statistically stationary and non-stationary features, (Hasselmann, 1976). We use a strict interpretation of “full statistically-nonstationary” to imply that “all” metrics of the simulation are converged. In our simulation only some statistics achieve stationarity during spinup, while other do not. The generative AI approach implicitly assumes stationarity of all statistics if it is to extrapolate future states from past ones. We will therefore note when the inference problem is applied to stationary versus non-stationary statistics.
Using Oceananigans.jl (Ramadhan et al., 2020; Silvestri et al., 2024; Wagner et al., 2025a; Silvestri et al., 2025; Wagner et al., 2025b), we solve the Boussinesq equations under the hydrostatic approximation. The dynamical core uses a nonlinear free surface and $z^{\star}$ coordinates. The prognostic variables are the horizontal velocities, $U$ and $V$, the SSH $\eta$, and the temperature $T$. A linear equation of state relates the buoyancy of seawater $B$ to the temperature, (Roquet et al., 2015). The figures will refer to the Oceananigans simulation output as “OcS” for brevity.
The idealized ocean domain extends from 15 to 75 degrees north in latitude and from 0 to 60 degrees in longitude with walls on all sides. The domain is 1800 m deep with a flat bottom. A zonal wind stress is applied at the surface, independent of longitude, with a cosine profile peaking eastward (positive) in the middle of the domain at a value of $0.1\text{ N}\text{m}^{-2}$ and at a minimum and westward (i.e. negative) at the northern and southern walls. The surface temperature is restored to a linear profile in latitude ranging from $30$ C on the southern boundary at $15^{\circ}$N, to 0 C at the northern boundary at $75^{\circ}$N. A quadratic bottom drag is imposed at the seafloor and convective adjustment is used for vertical mixing. The (dimensionless) quadratic drag coefficient is $10^{-3}$ and we use a horizontal diffusivity and viscosity of $2.5\times 10^{2}\text{ m}^{2}\text{s}^{-1}$, a background vertical diffusivity of $10^{-5}\text{ m}^{2}\text{s}^{-1}$, and background vertical viscosity of $10^{-2}\text{ m}^{2}\text{s}^{-1}$.
@@
Our simulation is not in statistical equilibrium for all observables of the flow, as can be observed from Figure 2, which shows the $L^{1}$ norm (the average absolute value) of all variables at a few selected depths, e.g.
@@
We use this measure as it is sensitive to localized fluctuations, such as the root-mean-squared value. Including metric terms does not change the overall trend and is unnecessary for the present case, but we will include them later when calculating physically meaningful quantities rather than statistical measures of non-stationarity. The initial spinup period (blue) is discarded. Thus, all the data in the work uses the time periods associated with the middle portion (red) and the end portion (orange) of the simulation. Closer to the surface (first column), it appears that a quasi-statistical equilibrium has been reached, whereas at depth (last column), statistical equilibrium has not been reached, especially for temperature. This quasi-equilibrium nature of the ocean is a feature we wished to capture in our dataset when investigating the limits of statistical inference predicated on past data. We comment that the sea surface height’s $L^{1}$ norm is also not in equilibrium. Stated differently, our “data distribution” (red) will be different than our “validation distribution” (orange) and will contribute as an extra source of error. That being said, there are many statistics which can, for the most part, be regarded as being in statistical equilibrium such as the surface horizontal velocity components.
@@
2.2 Probabilistic Models
In the present work, we draw samples from subsurface fields, $U,V,W,T$, given the sea surface height $\eta$, where the underlying probability distribution comes from the data distribution generated by the time series of the primitive equations. Mathematically, this is represented as a conditional probability distribution, $\mathcal{P}$,
$\text{conditional distribution}=\mathcal{P}(\text{subsurface fields}|\text{surface fields}).$ (2)
The conditional probability distribution outputs four fields $U,V,W,T$ at every grid point in three-dimensional space, conditioned on the SSH $\eta$. Concretely, samples from this distribution produce $4\times 256\times 256\times 15$ numbers conditioned on $256\times 256$ numbers using the full resolution of the simulation. This high-dimensional conditional probability distribution is learned from the data set using deep-learning methods, see Section 22.3.
If the conditional information is perfectly informative, there is a functional relationship, $\mathcal{F}$, between SSH and predicted fields and predicted fields. In this case, the conditional distribution collapses to a delta function,
$\mathcal{P}(\text{subsurface fields}|\text{surface fields})=\delta(\text{subsurface fields}-\mathcal{F}[{\text{surface fields}}]).$ (3)
If the provided information is “not useful” or maximally uninformative, then the conditional probability distribution, $\mathcal{P}$, should reduce down to the marginal distribution of the data, e.g., a random sample with statistics similar to those the generative model has been trained on independently of the surface information,
where $\mathcal{P}_{M}$ is the marginal distribution of the subsurface fields. Most inference tasks lie between these two extremes, Equations 3 and 4. We comment that we use the word “marginal” as is commonly used in the statistics literature with regard to distributions.
Traditional machine learning methods find a functional form $\mathcal{F}$ between input and outputs. This methodology is well-suited for problems where such a functional form exists. Otherwise, traditional methods that minimize a root-mean square (RMS) error to find a functional form can only infer the average value of the conditional probability in Equation 2, that is
$\mathcal{F}[\text{surface fields}]=\mathbb{E}_{\text{subsurface fields}}[\mathcal{P}(\text{subsurface fields}|\text{surface fields})],$ (5)
much like a “mean” statistic minimizes the RMS error of a random variable. (See Appendix A for a proof of this result.) If the conditional distribution is a delta-function, as in Equation 3, little is lost by performing the average; however, if there is significant variance in the output field, then the best that a deterministic model can do is to average over such variance. At the extreme, the output of a deterministic model will produce the ensemble average of the previously seen data, e.g. the average value of the marginal distribution, Equation 4. We surmise that this observation partially explains why fields drawn from a conditional distribution better represent the statistical properties of fluid flows compared with those learned via a deterministic functional relationship. Details will not be “washed away” from averaging.
An important consequence of Equation 5 is that we can compare generative models to traditional regression techniques indirectly by taking the ensemble average of the output given by the generative model. In this way we can see what details are missed by using non-generative approaches as well as provide a baseline for when we expect traditional approaches to outperform generative methods. This observation is independent of the neural network architecture and comes instead directly from the objective function that is being minimized by an “RMS loss” of a traditional method versus the “denoising score-matching loss” of a generative approach.
@@ 3.1 Baseline Performance of Generative Model
We see an overall agreement between the ground truth (second column) and the expected value of the model predictions (third column). This result means that a substantial ageostrophic component is also captured. The ensemble mean of the generated samples looks similar to the ground truth and the representative generated samples since there is little diversity in the generated samples. The conditional distribution is nearly a delta function, suggesting that the sea surface height is very informative of the ocean state and that a functional relationship can be found between the sea surface height and sub-surface currents. In other words, the present inference task is leaning towards Equation 3. Given that a traditional supervised learning task would produce the ensemble mean prediction of a generative model, low sample diversity would imply that a supervised learning task would perform well for the horizontal velocity components and the temperature field; however, the vertical velocity field has significant sample diversity (strong fluctuations in the northern region as can be seen by comparing the ensemble mean to the generated samples), implying that a traditional machine learning model would miss these fluctuations and “average them out”.
The sixth column shows the absolute value of the difference between the OcS and the AI mean, and the seventh column shows the ensemble standard deviation of the AI method. We represent predictive uncertainty by using the generated fields’ standard deviation at each grid point. The largest standard deviation is associated with the northern convecting region, especially for vertical velocity. We surmise that the sea surface height alone is insufficient to constrain the pointwise vertical velocity magnitude in these regions. A traditional machine learning model would average out these fluctuations to produce the average vertical velocity field as given by the “AI Mean” column. We see the flexibility of the generative model in informing our ability to apply traditional machine learning models. The second largest source of uncertainty is the eddying gyre region, which contributes to the uncertainty of the velocity components. The mean discrepancy column and the AI Std column have similar spatial structures, suggesting that the method can identify regions in which it is less capable of prediction.
Although mean discrepancy and the standard deviations are similiar in structure and magnitude across the velocity fields, we see that the temperature field has a different spatial structure. Although the strongly eddying region and the northern convective remain the largest source of uncertainty there is also a uniform background uncertainty in the domain as well. This uniform uncertainty of the temperature field could perhaps be related to the non-stationarity of the domain average temperature field, which is a statistic that evolves on a much longer timescale than others in the OcS. Ultimately, the relative errors of the temperature field are smaller as compared to the velocity fields. The standard deviation of the generative model’s temperature field overestimates this (small) uncertainty.
@@ 3.2 Ocean State Predictions and Uncertainties with Degraded Surface Height Input
These observations suggest that an SSH coarse-grained beyond $\mathcal{O}(1,000)$ kilometers is no longer informative since the generated samples become so diverse as to reproduce statistics of the data distribution, a “worst case” scenario for the generative method. As we coarse-grain the free surface (lose information), the inference task moves from something informative (similar to Equation 3) towards the uninformative (similar to Equation 4) marginal distribution of the training data. Taken together we see that the generative model can quantify the “non-existance” of a unique purely deterministic method for predicting subsurface states; the problem becomes inherently ill-posed from a deterministic perspective (i.e. trying to predict a single unique subsurface state is not possible, because the given information is insufficient to constrain potential subsurface states). For a deterministic method to succeed it must be supplemented with either additional information, some regularization or a criteria to select from the multitude of possible subsurface configurations, (e.g. choosing the most probable subsurface state, an ensemble mean subsurface state, or some other selection criteria).
@@ 3.3 Heat Flux Predictions and Uncertainties with Degraded Surface Height Input
$\overline{\cdot}\stackrel{{\scriptstyle\textrm{def}}}{{=}}\frac{1}{60^{\circ}}\int_{0}^{60^{\circ}}$ $\cdot d\lambda,$ (6)
@@
We also calculate the eddy heat fluxes as a function of SSH resolution at a $400$ meter depth in Figure 8. Similar to the previous plot, the top row is the vertical heat flux, the bottom row is the meridional heat flux, the “ground truth” results are the blue lines, and the generative AI results with 100 samples are shown as a ribbon with $0.6,0.7,0.8,0.9$ quantiles from darker to lighter shadings of red. The heat flux prediction grows in uncertainty as we coarse-grain the sea surface height. The loss of information here is understood upon examining the first column of Figure 4. As the sea surface height (SSH) field is increasingly coarse-grained, we lose precise information about the location of the jet separation point — the region where the western boundary current detaches from the wall and begins to meander eastward. This separation point is critical in organizing mesoscale eddy activity and lateral heat transport. When spatial detail in the SSH is removed, the model can no longer resolve where the jet begins to meander, leading to increased uncertainty in the predicted eddy heat fluxes, particularly in the downstream region where the current interacts with ambient stratification. The ability of the generative model to produce plausible eddy fluxes in the ill-posed regime is another feature of generative models: a deterministic model would be relegated to making an approximation of the form $\mathbb{E}[vT]\approx\mathbb{E}[v]\mathbb{E}[T]$ where the expected value on the right hand-side is the output of the deterministic model’s prediction for $v$ and $T$. This occurs for the same reason mentioned prior that a traditional machine learning model can (at best) produce the ensemble mean of what a generative model can produce.